# compat=1.0.0

### v1

@@ Abstract
We investigate the thermal production of charm quark in the strongly interacting quark-gluon plasma (sQGP) created in heavy-ion collisions at relativistic energies. Our study is based on the off-shell parton-hadron-string dynamics (PHSD) transport approach describing the full time evolution of heavy-ion collisions on a microscopic bases with hadronic and partonic degrees-of-freedom. The sQGP is realized within the effective dynamical quasi-particle model (DQPM) which is adjusted to reproduce the lattice QCD results for the thermodynamic observables of the sQGP. Relying that the DQPM successfully describes the spatial diffusion coefficients $D_{s}$ from the lQCD, which controls the interaction of charm quarks with thermal partons (expressed in terms of strongly interacting off-shell quasiparticles), we evaluate the production of pairs of anti-charm/charm quarks through the rotation of Feynman diagrams using cross symmetry for elastic scattering of charm quark with thermal partons ((anti-)quarks or gluons) in the sQGP by the exchange of the incoming charm quark and outgoing light parton in elastic scattering diagrams. The charm quark annihilation is realized by detailed balance. We find that the number of produced thermal charm quark pairs strongly depends on the charm quark mass in the QGP. While for the heavy charm quarks of mass $m_{c}=1.8$ GeV it is subdominant compared to the primary charm production by binary nucleon-nucleon collisions at RHIC and LHC energies, the numbers of primary and thermal charm quarks become comparable for a smaller (bare) $m_{c}=1.2$ GeV. Comparing with the experimental data on the $R_{\rm AA}$ of $D$-mesons in heavy-ion collisions at RHIC and LHC energies, it is more favorable for charm quarks in the QGP to gain additional mass due to thermal effects rather than to have a low bare mass.
@@
Heavy flavors (charm and beauty quarks/antiquarks) are one of the promising tomographic probes of the properties of hot and dense matter created in relativistic heavy-ion collisions. There are extended experimental Adamczyk et al. (2014); Adam et al. (2019); Sirunyan et al. (2018); Acharya et al. (2022) and theoretical efforts (cf. e.g. Refs.  Gossiaux et al. (2009); He et al. (2012); Uphoff et al. (2012); Cao et al. (2013); Song et al. (2015); Plumari et al. (2018); Cao et al. (2020); Beraudo et al. (2022) as well as the reviews Cao et al. (2019); Beraudo et al. (2018); Xu et al. (2019); Zhao et al. (2023)) in the last decade to improve our understanding of the production mechanisms of heavy flavours and their interactions in the quark-gluon plasma (QGP) and in the hadronic corona after hadronization. Since heavy flavors are massive, they dominantly produced by primary nucleon-nucleon scatterings in heavy-ion collisions. Therefore, the number of produced heavy flavors scales with the number of nucleon binary collisions. However, heavy flavors can additionally be produced by partonic reactions in the QGP - e.g. quark-antiquark annihilation or gluon-gluon fusion, which we denote by ”thermal” production mechanism in the QGP even if it happens at early reaction time, when the QGP doesn’t equilibrate yet. With increasing collision energy, the temperature of produced QGP grows, and the possibility for a thermal production of heavy flavor increases. This goes along in competition with the backward reactions of charm-anticharm quark annihilation. If the system achieves thermal equilibrium, the forward and backward reactions compensate each other and the fugacity (which control the deviation from a statistical phase space distribution) of the charm quarks will approach 1 Mykhaylova (2022). However, in realistic heavy-ion collisions the system is far from equilibrium in the initial phase and also deviates from equilibrium during the fast expansion phase not only chemically - as shown in the Statistical Hadronization Model (SHM) Andronic et al. (2021) - but also kinetically. Thus, one expects a violation of the balance between production/annihilation of charm-anticharm quarks produced in the QGP during the heavy-ion collisions.
@@
The interactions of heavy flavor with thermal partons in the QGP are partly known from the spatial diffusion coefficients of heavy flavor which are provided by lQCD Banerjee et al. (2012); Altenkort et al. (2024). The properties of a thermalized strongly interacting QGP has been extensively studied with the Dynamical-QuasiParticle Model (DQPM) Peshier and Cassing (2005); Cassing (2007a, b); Berrehrah et al. (2016a); Moreau et al. (2019); Soloveva et al. (2021). The DQPM is an effective model that provides a microscopic picture of the thermalized QGP, reprodusing its macroscopic properties, such as the equation of state from lQCD, in terms of quasiparticles whose pole masses and spectral widths depend on temperature and baryon chemical potential Moreau et al. (2019). The DQPM is applied to parton interactions by using the leading-order Feynman diagrams and reproduces the spatial diffusion coefficient of heavy quark from lQCD in a wide range of temperatures Berrehrah et al. (2014a); Song et al. (2020a). Based on this success, we extend the calculations to charm quark production in the QGP and in heavy-ion collisions by changing the incoming charm quark to the outgoing anticharm quark in the Feynman diagrams for the elastic scattering of a charm quark.
@@
Since the charm quark mass is higher than typical temperatures of the matter produced in heavy-ion collisions, the thermal production of charm quarks is usually ignored. However, the initial temperature of the sQGP, especially at the top energy of LHC, is considerably high. Furthermore, the thermal masses of quarks and gluons increase with temperature in the DQPM, which is supported by the thermal quantum field theory  Kapusta and Gale (2011); Bellac (2011). Both effects enhance the possibility for charm quarks to be thermally produced in heavy-ion collisions.
@@
The goal of our study is to estimate the amount of thermal charm-anticharm quark production by partonic interactions in the sQGP in addition to the primary charm/anticharm production by initial binary nucleon-nucleon reactions in central heavy-ion collisions at RHIC and LHC and investigate their properties in comparison with the experimental data. Furthermore, we explore the balance between charm-anticharm quark production and annihilation in the expanding QGP matter created during the heavy-ion collisions.
@@
Assuming local thermal equilibrium, one can rely on hydrodynamics for the time-evolution of nuclear matter in heavy-ion collisions. However, the initial thermalization time, from which hydrodynamics is applicable, is uncertain. On the other hand, this initial time is most important for the ”thermal”/QGP production of charm due to its high temperature and/or high energy density. Moreover, in the initial phase of heavy-ion collisions the system is out-of equilibrium such that an application of hydrodynamical models at very early times is not possible.
@@
In order to overcome thse difficulties, we base our study of charm dynamics on a nonequilibrium transport approach, the parton-hadron-string dynamics (PHSD), which is a microscopic transport approach based on first order gradient expansion of the Kadanoff-Baym equations taking into account the off-shellness of particles for their strong interactions Cassing and Bratkovskaya (2008, 2009); Cassing (2009); Bratkovskaya et al. (2011); Linnyk et al. (2016). Based on the DQPM for the description of the sQGP phase, it describes the experimental data on open and hidden heavy flavors in relativistic heavy-ion collisions Song et al. (2015, 2016, 2017).
@@
This paper is organized as follows: In Sec. II we briefly recall the DQPM. In Sec. III the heavy quark production in the DQPM and the thermal production rate of charm quarks in the QGP are introduced. In Sec. IV we study the thermal charm production in heavy-ion collisions including charm quark pair annihilation by the detailed balance. The dependence on the charm quark thermal mass is investigated too in the comparison with experimental data. We, furthermore, study the dependence of the charm diffusion coefficient $D_{s}$ on the choice of the charm quark mass $m_{c}$. Finally, a summary is given in Sec. V.
@@ II Dynamical Quasiparticle Model
Here we briefly recall the basic ideas of the Dynamical Quasiparticle Model (DQPM) Peshier and Cassing (2005); Cassing (2007a, b); Berrehrah et al. (2016a); Moreau et al. (2019); Soloveva et al. (2021) which is an effective model for the description of the sQGP in terms of strongly interacting quasiparticles (quarks and gluons) matched to reproduce the results of lattice QCD calculations in thermal equilibrium and at vanishing chemical potential. The quasiparticles are characterized by single-particle (two-point) Green’s functions, i.e. ”dressed” propagators:
@@
for quarks, antiquarks, and gluons ($j=q,\bar{q},g$), using $\omega=p_{0}$ for energy, the widths $\gamma_{j}$ and the masses $M_{j}$ and the complex self-energies for gluons $\Pi=M_{g}^{2}-2i\omega\gamma_{g}$ and for (anti)quarks $\Sigma_{q}=M_{q}^{2}-2i\omega\gamma_{q}$, where the real part of the self-energies is associated with dynamically generated thermal masses, while the imaginary part provides information about the lifetime and reaction rates of the particles.
We use an ansatz for the pole masses $M_{j}(T,\mu_{q})$ and widths $\gamma_{j}(T,\mu_{q})$ as functions of the temperature $T$ and the quark chemical potential $\mu_{q}$ based on the HTL thermal mass in the asymptotic high-temperature regime Bellac (2011); Linnyk et al. (2016):
@@
where $N_{c}\ (=3)$ stands for the number of colors and $N_{f}\ (=3)$ denotes the number of light flavors.
@@
Here $c_{m}=14.4$ is related to a magnetic cutoff, which is an additional parameter in the DQPM, while $C_{q}=\dfrac{N_{c}^{2}-1}{2N_{c}}=4/3$ and $C_{g}=N_{c}=3$ are the QCD color factors for quarks and gluons, respectively. We also assume that all (anti)quarks have the same $T$-dependence for the width.
The thermal properties of quasiparticles and their interactions (defined via transport coefficients) strongly depend on the coupling constant $g$ which accounts for nonperturbative effects and enters the definition of the masses and widths of quasiparticles – Eqs. (3), (4). In the DQPM $g^{2}$ is extracted from lQCD data on the entropy density $s$ by a parametrization method introduced in Ref. Berrehrah et al. (2016b) using the scaling of the ratio $s(T,g^{2})/T^{3}$ versus $T$ for a given value of $g^{2}$:
@@
Here $d=169.934,e=-0.178434$ and $f=1.14631$ are the dimensionless parameters obtained by adjusting the quasiparticle entropy density $s(T,\mu_{B}=0)$ to the lQCD data provided by the BMW Collaboration Borsanyi et al. (2012, 2014). $s^{QCD}_{SB}=19/9\pi^{2}T^{3}$ is the Stefan-Boltzmann entropy density. The extension of the coupling constant to finite baryon chemical potential $\mu_{B}$ is realized using a scaling hypothesis Cassing (2009) that works up to $\mu_{B}\approx 500$ MeV.
@@ III Thermal production of charm quarks in the QGP
The thermal production of charm quarks in a QGP is closely related to the elastic scattering of charm quarks in the QGP. Figs. 1 and 2 show the Feynman diagrams for thermal production from quark and antiquark annihilation and two gluon fusion, respectively. They are obtained by rotation of Feynman diagrams for charm quark elastic scattering of light (anti)quarks or gluons in Ref. Berrehrah et al. (2014b). The calculations of the Feynman diagrams are realized by exchanging the incoming charm quark and the outgoing light quark or a gluon in the elastic scattering diagrams. For example, the scattering amplitudes squared are related to each other as follows:
@@
We note that Figs. 1 and 2 are not simple leading-order diagrams in the DQPM, because the exchanged quark and gluon have a dressed mass and width from the interactions in the QGP. The mass and width depend on temperature and baryon chemical potential and are extracted from the equation of state in lQCD Moreau et al. (2019).
Since charm quark elastic scattering in the DQPM is quantitatively supported by lQCD in terms of the spatial diffusion coefficients Song et al. (2020a), the thermal production of charm quarks in the DQPM is expected to be reliable.
Refer to captionRefer to caption
Figure 3: (Top) Scattering cross sections of a light quark-antiquark pair and of two gluons to produce a charm quark pair as a function of scattering energy $\sqrt{s}$ at $T=T_{c}$, 2 $T_{c}$ and 3 $T_{c}$ and (bottom) the production rate of charm quark pairs in the QGP as a function of temperature. The charm quark mass is taken to be 1.5 GeV
Fig. 3 shows scattering cross sections for charm production from quark-antiquark annihilation and two gluon fusion as a function of scattering energy at the temperatures of 1, 2, and 3 $T_{c}$. The charm quark mass is taken to be 1.5 GeV. Both cross sections increase rapidly above the threshold energy and then decrease. Since the strong coupling has a maximum at $T_{c}$ and decreases with increasing temperature, the production cross sections are largest at $T_{c}$ and much smaller at 2 and 3 $T_{c}$. The lower panel shows the production rate of charm quark pairs per unit time and unit volume is defined by
@@
×$\times\int\frac{d^{3}k_{1}d^{3}k_{2}}{(2\pi)^{6}}f_{g}(k_{1},m_{1})f_{g}(k_{2},m_{2})v_{gg}\sigma_{gg\rightarrow c\bar{c}},$ (9)
where $\rho_{i}$ and $f_{i}$ are the spectral functions and momentum distribution function of parton $i$ at temperature $T$, respectively; $v_{ij}$ is the relative velocity of partons $i$ and $j$, while $\sigma$ is the scattering cross section.
Although the production cross section has a maximum at $T_{c}$, the production rate is quite small, because only a few reactions of quark-antiquark annihilation or of two gluon fusion are above the threshold energy for charm pair production at low temperature. Comparing the two channels, the quark-antiquark annihilation has a slightly larger production rate than two gluon fusion, although the difference is small.

### v4

@@ Abstract
We investigate the thermal production of charm quarks in the strongly interacting quark-gluon plasma (sQGP) created in heavy-ion collisions at relativistic energies. Our study is based on the off-shell parton-hadron-string dynamics (PHSD) transport approach describing the full time evolution of heavy-ion collisions on a microscopic basis with hadronic and partonic degrees of freedom. The sQGP is realized within the effective dynamical quasi-particle model (DQPM) which is adjusted to reproduce the lattice QCD results for the thermodynamic observables of the sQGP. Relying on the fact that the DQPM successfully describes the spatial diffusion coefficients $D_{s}$ from the lQCD, which control the interaction of charm quarks with thermal partons (expressed in terms of strongly interacting off-shell quasiparticles), we evaluate the production of charm quark pairs through the rotation of Feynman diagrams such that the incoming charm quark and outgoing light parton in elastic scattering diagrams are exchanged. The charm quark annihilation is realized by detailed balance. We find that the number of produced thermal charm quark pairs strongly depends on the charm quark mass in the QGP. While for the heavy charm quarks of mass $m_{c}=1.8$ GeV it is subdominant compared to the primary charm production by binary nucleon-nucleon collisions at RHIC and LHC energies, the numbers of primary and thermal charm quarks become comparable for a smaller (bare) $m_{c}=1.2$ GeV. Compared with the experimental data on the $R_{\rm AA}$ of $D$-mesons in heavy-ion collisions at RHIC and LHC energies, it is more favorable for charm quarks in the QGP to gain additional mass due to thermal effects rather than to have a low bare mass.
@@ I Introduction
Heavy flavors (charm and beauty quarks/antiquarks) are one of the promising tomographic probes of the properties of hot and dense matter created in relativistic heavy-ion collisions. There have been extended experimental Adamczyk et al. (2014); Adam et al. (2019); Sirunyan et al. (2018); Acharya et al. (2022) and theoretical efforts (cf. Refs.  Gossiaux et al. (2009); He et al. (2012); Uphoff et al. (2012); Cao et al. (2013); Song et al. (2015); Plumari et al. (2018); Cao et al. (2020); Beraudo et al. (2022) as well as the reviews Cao et al. (2019); Beraudo et al. (2018); Xu et al. (2019); Zhao et al. (2023)) in the last decade to improve our understanding of the production mechanisms of heavy flavours and their interactions in the quark-gluon plasma (QGP) and in the hadronic corona after hadronization. Since heavy flavors are massive, they are dominantly produced by primary nucleon-nucleon scatterings in heavy-ion collisions. Therefore, the number of produced heavy flavors scales with the number of nucleon binary collisions. However, heavy flavors can additionally be produced by partonic reactions in the QGP – e.g., quark-antiquark annihilation or gluon-gluon fusion, which we denote by ”thermal” production mechanism in the QGP even if it may happen at early reaction time, when the QGP does not equilibrate yet. With increasing collision energy, the temperature of produced QGP grows, and the possibility for thermal production of heavy flavor increases. This goes along in competition with the backward reactions of charm-anticharm quark annihilation. If the system achieves chemical equilibrium, the forward and backward reactions compensate each other, and the fugacity (which controls the deviation from a statistical phase-space distribution) of the charm quarks will approach 1 Mykhaylova (2022). However, in realistic heavy-ion collisions, the system is far from equilibrium in the initial phase and also deviates from equilibrium during the fast expansion phase, not only chemically – as shown in the Statistical Hadronization Model (SHM) Andronic et al. (2021) – but also kinetically. Thus, one expects a violation of the balance between production/annihilation of charm-anticharm quarks produced in the QGP during the heavy-ion collisions.
Thermal production has been studied for more than three decades. Because charm production requires energetic scattering of partons, it was found that the early stage of heavy-ion collisions is important Muller and Wang (1992); Geiger (1993); Levai et al. (1995), and massive partons are more effective for it Levai and Vogt (1997). In pQCD, calculations have been carried out up to next-to-leading order Zhang et al. (2008). These calculations commonly predicted that the thermal charm production would have sizable effects in heavy-ion collisions at high energies, such as those at the LHC. However, charm production, even at the LHC, is well explained in the SHM together with shadowing effects, which suppress charm production at midrapidity and at small $p_{T}$, corresponding to small $x$ in the parton distribution function. To understand this success, it is necessary to use reliable cross sections for charm production in parton scattering.
@@
The interactions of heavy flavor with thermal partons in the QGP are partly known from the spatial diffusion coefficients of heavy flavor, which are provided by lQCD Banerjee et al. (2012); Altenkort et al. (2024). The properties of a thermalized, strongly interacting QGP have been extensively studied with the Dynamical-QuasiParticle Model (DQPM) Peshier and Cassing (2005); Cassing (2007a, b); Berrehrah et al. (2016a); Moreau et al. (2019); Soloveva et al. (2021). The DQPM is an effective model that provides a microscopic picture of the thermalized QGP, reproducing its macroscopic properties, such as the equation of state (EoS) from lQCD, in terms of quasiparticles whose pole masses and spectral widths depend on temperature and baryon chemical potential Moreau et al. (2019). The DQPM is applied to parton interactions by using the leading-order Feynman diagrams and reproduces the spatial diffusion coefficient of heavy quark from lQCD in a wide range of temperatures Berrehrah et al. (2014a); Song et al. (2020a).
We stress that since the DQPM is based on propagator 2-particle irreducible representation, it is free from the pQCD divergences and necessity to introduce Debye masses for the regularisation of scattering amplitudes. Indeed, the properties of quasiparticles expressed in terms of complex self energies – with the real part related to the thermal paron masses and imaginary part to their widths – are matched to reproduce the thermodynamic properties of lQCD at $\mu_{B}=0$ and extrapolates to finite $\mu_{B}$ using a scaling hypothesis. Thus, the DQPM propagators effectively includes the resummation effects since the self-energies are fitted to the lQCD data which a priori contain the full QCD solutions.
Based on this success, we extend the calculations to charm quark production in the QGP and in heavy-ion collisions by changing the incoming charm quark to the outgoing anticharm quark in the Feynman diagrams for the elastic scattering of a charm quark.
@@
Since the charm quark mass is higher than the typical temperatures of the matter produced in heavy-ion collisions, the thermal production of charm quarks is usually ignored. However, the initial temperature of the sQGP, especially at the top energy of the LHC, is considerably high. Furthermore, the thermal masses of quarks and gluons increase with temperature in the DQPM, which is supported by the thermal quantum field theory  Kapusta and Gale (2011); Bellac (2011). Both effects enhance the possibility of charm quarks being thermally produced in heavy-ion collisions.
@@
The goal of our study is to estimate the amount of thermal charm-anticharm quark production by partonic interactions in the sQGP, in addition to the primary charm/anticharm production by initial binary nucleon-nucleon reactions in central heavy-ion collisions at RHIC and the LHC, and investigate their properties in comparison with the experimental data. Furthermore, we explore the balance between charm-anticharm quark production and annihilation in the expanding QGP matter created during heavy-ion collisions.
@@
Assuming local thermal equilibrium, one can rely on hydrodynamics for the time evolution of nuclear matter in heavy-ion collisions. However, the initial thermalization time, from which hydrodynamics is applicable, is uncertain. On the other hand, this initial time is most important for the thermal/QGP production of charm due to its high temperature and/or high energy density. Moreover, in the initial phase of heavy-ion collisions, the system is out of equilibrium, so the application of hydrodynamical models at very early times is not possible.
@@
In order to overcome these difficulties, we base our study of charm dynamics on a nonequilibrium transport approach, the parton-hadron-string dynamics (PHSD), which is a microscopic transport approach based on the first order gradient expansion of the Kadanoff-Baym equations, taking into account the off-shellness of particles for their strong interactions Cassing and Bratkovskaya (2008, 2009); Cassing (2009); Bratkovskaya et al. (2011); Linnyk et al. (2016). Based on the DQPM for the description of the sQGP phase, it describes the experimental data on open and hidden heavy flavors in relativistic heavy-ion collisions Song et al. (2015, 2016, 2017).
@@
This paper is organized as follows: in Sec. II, we briefly recall the DQPM. In Sec. III, the heavy quark production in the DQPM and the thermal production rate of charm quarks in the QGP are introduced. In Sec. IV, we study the thermal charm production in heavy-ion collisions including charm quark pair annihilation by the detailed balance. The dependence on the charm quark thermal mass is also investigated in comparison with experimental data. Furthermore, we study the dependence of the charm diffusion coefficient $D_{s}$ on the choice of the charm quark mass $m_{c}$. Finally, a summary is given in Sec. V.
@@ II Dynamical Quasiparticle Model
Here, we briefly recall the basic ideas of the Dynamical Quasiparticle Model (DQPM) Peshier and Cassing (2005); Cassing (2007a, b); Berrehrah et al. (2016a); Moreau et al. (2019); Soloveva et al. (2021), which is an effective model for the description of the sQGP in terms of strongly interacting quasiparticles (quarks and gluons) matched to reproduce the results of lattice QCD calculations in thermal equilibrium and at vanishing chemical potential. The quasiparticles are characterized by single-particle (two-point) Green’s functions, i.e., ”dressed” propagators:
@@
for quarks, antiquarks, and gluons ($j=q,\bar{q},g$), using $\omega=p_{0}$ for energy, the widths $\gamma_{j}$, the masses $M_{j}$, and the complex self-energies for gluons $\Pi=M_{g}^{2}-2i\omega\gamma_{g}$ and for (anti)quarks $\Sigma_{q}=M_{q}^{2}-2i\omega\gamma_{q}$, where the real part of the self-energies is associated with dynamically generated thermal masses, while the imaginary part provides information about the lifetime and reaction rates of the particles.
We use an ansatz for the pole masses $M_{j}(T,\mu_{q})$ and widths $\gamma_{j}(T,\mu_{q})$ as functions of the temperature $T$ and the quark chemical potential $\mu_{q}$, based on the HTL thermal mass in the asymptotic high-temperature regime Bellac (2011); Linnyk et al. (2016):
@@
where $N_{c}\ (=3)$ stands for the number of colors, and $N_{f}\ (=3)$ denotes the number of light flavors.
@@
Here, $c_{m}=14.4$ is related to a magnetic cutoff, which is an additional parameter in the DQPM, while $C_{q}=\dfrac{N_{c}^{2}-1}{2N_{c}}=4/3$ and $C_{g}=N_{c}=3$ are the QCD color factors for quarks and gluons, respectively. We also assume that all (anti)quarks have the same $T$ dependence for the width.
The thermal properties of quasiparticles and their interactions (defined via transport coefficients) strongly depend on the coupling constant $g$, which accounts for nonperturbative effects and enters the definition of the masses and widths of quasiparticles – Eqs. (3), (4). In the DQPM, $g^{2}$ is extracted from lQCD data on the entropy density $s$ by a parametrization method introduced in Ref. Berrehrah et al. (2016b), using the scaling of the ratio $s(T,g^{2})/T^{3}$ versus $T$ for a given value of $g^{2}$:
@@
Here $d=169.934,e=-0.178434$, and $f=1.14631$ are the dimensionless parameters obtained by adjusting the quasiparticle entropy density $s(T,\mu_{B}=0)$ to the lQCD data provided by the BMW Collaboration Borsanyi et al. (2012, 2014), and $s^{QCD}_{SB}=19/9\pi^{2}T^{3}$ is the Stefan-Boltzmann entropy density. The extension of the coupling constant to finite baryon chemical potential $\mu_{B}$ is realized using a scaling hypothesis Cassing (2009) that works up to $\mu_{B}\approx 500$ MeV.
Figure 1: Strong coupling $\alpha_{s}$ as a function of scaled temperature at $\mu_{B}=$ 0, 0.2, and 0.4 GeV, compared with lattice results.
As shown in Fig. 1, the resulting $\alpha_{s}$ starts with a small value at high temperature but increases with decreasing temperature and reaches 2.9 for $\mu_{B}=0$ and 2.0 for $\mu_{B}=0.6$ GeV at $T_{c}$ Moreau et al. (2019); Grishmanovskii et al. (2024), which is consistent with the lattice QCD results Kaczmarek et al. (2004); Kaczmarek and Zantow (2005); Kaczmarek (2007).
@@ III Thermal production of charm quarks in the QGP
The thermal production of charm quarks in a QGP is closely related to the elastic scattering of charm quarks in the QGP. Figures 2 and 3 show the Feynman diagrams for thermal production from quark-antiquark annihilation and two-gluon fusion, respectively. They are obtained by rotating the Feynman diagrams for charm quark elastic scattering with light (anti)quarks or gluons in Ref. Berrehrah et al. (2014b). The calculations of the Feynman diagrams are performed by exchanging the incoming charm quark with the outgoing light quark or a gluon in the elastic scattering diagrams. For example, the squared scattering amplitudes are related to each other as follows:
@@
We note that Figs. 2 and 3 are not simple leading-order diagrams in the DQPM, as the exchanged quark and gluon have dressed masses and widths from the interactions in the QGP. The mass and width depend on temperature and baryon chemical potential, and are extracted from the lQCD equation of state Moreau et al. (2019).
Since charm quark elastic scattering in the DQPM is quantitatively supported by lQCD in terms of the spatial diffusion coefficients Song et al. (2020a), the thermal production of charm quarks in the DQPM is expected to be reliable. In the case of elastic scattering the dressed mass and spectral width of light partons play important role to remove singularities and reduce forward scattering compared to the scattering of charm quark with massless partons Berrehrah et al. (2014b). However, this role is not manifested in the case of charm production, because $t-$channel in the elastic scattering turns to $s-$channel in the charm production, which does not have singularity, and the forward scattering is suppressed in the charm production even for the massless light partons.
Figure 4: (Top) scattering cross sections of a light quark-antiquark pair and of two gluons to produce a charm quark pair as a function of scattering energy $\sqrt{s}$ at $T=T_{c}$, 2 $T_{c}$, and 3 $T_{c}$; (bottom) the production rate of charm quark pairs in the QGP as a function of temperature. The charm quark mass is taken to be 1.5 GeV.
Figure 4 shows the scattering cross sections for charm production from quark-antiquark annihilation and two-gluon fusion as a function of scattering energy at the temperatures of 1, 2, and 3 $T_{c}$. The charm quark mass is taken to be 1.5 GeV. Both cross sections increase rapidly above the threshold energy and then decrease. Since the strong coupling has a maximum at $T_{c}$ and decreases with increasing temperature, the production cross sections are largest at $T_{c}$ and much smaller at 2 and 3 $T_{c}$. The lower panel shows the production rate of charm quark pairs per unit time and unit volume, defined by
@@
×$\times\int\frac{d^{3}k_{1}d^{3}k_{2}}{(2\pi)^{6}}f_{g}(k_{1},m_{1})f_{g}(k_{2},m_{2})v_{gg}\sigma_{gg\rightarrow c\bar{c}};$ (9)
where $\rho_{i}$ and $f_{i}$ are the spectral functions and momentum distribution function of parton $i$ at temperature $T$, respectively; $v_{ij}$ is the relative velocity of partons $i$ and $j$, and $\sigma$ is the scattering cross section.
Although the production cross section has a maximum at $T_{c}$, the production rate is quite small because only a few reactions of quark-antiquark annihilation or two-gluon fusion are above the threshold energy for charm pair production at low temperature. When comparing the two channels, quark-antiquark annihilation has a slightly larger production rate than two-gluon fusion, although the difference is small. We note that the off-shell effect of light partons enhances the charm production rate from gluon fusion by about 7%, while the enhancement of the rate from quark annihilation is less than 1%.
We also note that our $2\to 2$ production rates are in line with those from Ref. Zhang et al. (2008) which considers not only 2$\rightarrow$2 reactions but also next-to-leading order 2$\rightarrow$3 reactions ($q+\bar{q}\to c+\bar{c}+g$ and $g+g\to c+\bar{c}+g$) within the pQCD framework, which give a visible contribution to the production of charm pairs. However, for the case of non-perturbative QCD degrees-of-freedom of the DQPM, it has been shown in Ref. Grishmanovskii et al. (2024) that the contribution of the $2\to 3$ processes for the emission of a massive gluon to the thermal rate in the equilibrated QGP is suppressed contrary to the elastic $2\to 2$ scattering (cf. Fig. 16 of Ref. Grishmanovskii et al. (2024)). Thus, we expect that the processes $q+\bar{q}\to c+\bar{c}+g$ and $g+g\to c+\bar{c}+g$ would be even more strongly suppressed due to the large mass of emitted gluon in thermal equilibrium. To clarify this issue quantitatively would require a further study. Moreover, since in Ref. Zhang et al. (2008) the smaller charm quark mass is used and 2$\rightarrow$3 reactions are included, the obtained production rates are larger than in this study. However, the rates are applied in Ref. Zhang et al. (2008) only after the initial thermalization time of a fireball, which will be overcome in our study by using the production cross sections directly, as shown in the next section.